Marco got 25% on a test out of 80. How many marks out of 80 did he earn?
1
Expert's answer
2020-07-20T18:22:33-0400

Percentage means out of hundred ( per cent ).

So, if the total marks were hundred , then Marco would have got 25% marks, which is equal to 25 marks. Hence Marco is getting 1/4 th of the total marks ( 25/100 = 1/4 ).

Now when the actual full marks is 80, Marco gets 25% or 1/4 th of the full marks, that of 80;

1/4 th of 80 is equal to 20.

Hence , Marco gets 20 marks out of 80.


The above calculations can be summarised as follows :

25% { 25/100 } of 80 = (25/100) * 80 = (1/4) * 80 = 20.


Answer :

Marco earned 20 marks out of 80 in the test.